x^2/3-y^2/3 / x^1/3 + y^1/3
Answer by htmentor(1343)

If we make the substitution w = x^1/3, z = y^1/3, we can eliminate the fractional exponents
(x^2/3 - y^2/3)/(x^1/3 + y^1/3) -> (w^2 - z^2)/(w + z)
The numerator can be factored as (w+z)(w-z):
(w+z)(w-z)/(w+z) = w-z
Substituting back the original variables gives x^1/3 - y^1/3